5) $8000 is put into an empty savings account with a nominal interest rate (APR) of 1.5%. No other contributions are made to the account. With monthly compounding, how much interest will have been earned after five years? a) $625 b) $635 c) $645 d) $655
Answer #1

Answer 5

The correct answer is (a) $625

Formula;

With Monthly compounding Amount is given by:

A = P(1 + r/12)12n

Here A = Amount after n years , n = 5 , r = Annual interest rate = 1.5/100 = 0.015 amd P = Principle = 8000

=> A = 8000(1 + 0.015/12)12*5

=> A = 8625(Closest Value)

Hence, Interest Rate earned = A - P = 8625 - 8000 = 625(approx)

Hence, the correct answer is (a) $625